KCLO3 decomposes to KCL and O2.if the volume of O2 obtained in this reaction is 1.12 litres at STP,weigh of KCL formed in the reaction is

Grade:11

1 Answers

Lovedeep
4 years ago
Applying poac on the reaction
Kclo3----->kcl+o2
3*mol of kclo3=2*mol of o2-------eqn1;
1*mol of kclo3=1*mol of kcl------eqn2;
      By eqn 1 and 2
3*mol of kcl=2"mol of o2
 
And 22.4litre=1mol
Then 1.12lit=1/20mol
 
33*mol of kcl=2*(1/20)
Mol of kcl=(2/3)*(1/20)
X/74.5=(2/3)*(1/20)
X=2.48gm
 
3*mol of kcl